# Bondi-type accretion onto a Kerr black hole in the kinetic regime

### v2

@@ Abstract
I Introduction
The search for an exact solution representing a Bondi-type accretion in Kerr spacetime is a long-standing challenge. In 1988 Petrich, Shapiro, and Teukolsky obtained an analytic solution, assuming a zero-vorticity potential flow and the ultra-hard equation of state [1]. Their method was based on the observation that for the ultra-hard equation of state, the equations of motion governing zero-vorticity perfect fluids can be expressed in terms of a potential, and they are linear. More precisely, the continuity equation can be written as the massless Klein-Gordon equation, which is separable on the Kerr background spacetime. Unfortunately, no straightforward generalizations to perfect fluids with more realistic equations of state seem to be possible (see [2, 3, 4] for attempts to obtain approximate solutions).
@@ I Introduction
By a Bondi-type accretion we mean a nearly spherically symmetric, stationary flow of gas onto the black hole. The gas extends to infinity, where it is assumed to be homogeneous and at rest. Within the general-relativistic kinetic approximation, such models were investigated in [5, 6, 7, 8, 9, 10, 11], assuming spherical symmetry. Properties of other kinetic models in the Kerr spacetime were analyzed in [12, 13, 14, 15]. Numerical simulations of accretion flows of the collisionless magnetized gas onto Kerr black holes were performed in [16, 17, 18, 19].
@@
II Kinetic description of matter in the Kerr spacetime
Let $(\mathcal{M},g)$ and $T_{x}^{\ast}\mathcal{M}$ denote the spacetime manifold and the cotangent space at $x\in\mathcal{M}$, respectively. The gas is described in terms of the one-particle distribution function $f\colon\Gamma_{m}^{+}\to\mathbb{R}$ defined on the future mass shell $\Gamma_{m}^{+}:=\bigcup_{x\in\mathcal{M}}\{(x,p)\colon p\in P_{x}^{+}(m)\}$, where
@@ II Kinetic description of matter in the Kerr spacetime
$p\text{ is future-directed}\}.$
@@
Geodesic equations are separable [21, 22] in the coordinates $(t,r,\vartheta,\varphi)$ as follows: $p_{t}=E$, $p_{\varphi}=L_{z}$, $p_{\vartheta}^{2}=\Theta(\vartheta)$, $\left(\Delta p_{r}-2MEr+aL_{z}\right)^{2}=R(r)$, where $\Delta:=r^{2}-2Mr+a^{2}$,
@@
III Bondi-type solution
The components of the particle current density associated with absorbed and scattered orbits can be expressed as
@@ IV Accretion rates
$\lim\limits_{z\rightarrow\infty}\frac{1}{\sqrt{2\pi z}}\frac{\mathcal{\dot{J}}}{\varepsilon_{\infty}}$ $=$ $\alpha\left(\frac{32}{3}+\frac{2}{5}\alpha^{2}\right)M^{3},$ (20b)
[…]
$\lim\limits_{z\rightarrow 0}\frac{\mathcal{\dot{J}}}{\varepsilon_{\infty}}$ $=$ $36\pi\alpha M^{3}.$ (20d)
@@ V Discussion
up to third order in the rotation parameter, where
Equations (21) imply a decrease of the spin parameter magnitude at the rate
$\dot{\alpha}=-2\varepsilon_{\infty}\mathcal{A}(1+\delta)(1-\mathcal{C}\alpha^{2})M\alpha,\quad\mathcal{C}=\frac{\mathcal{B}-\delta\gamma}{1+\delta},$ (21c)
where we have used the relation $J=\alpha M^{2}$. Equations (21a) and (21c) govern the time evolution of the black hole mass and spin parameter, and they yield the following relation:
@@
The second scenario corresponds to a supermassive black hole which accretes dark matter from its environment. Assuming a reservoir of constant density and low temperature and integrating Eq. (23), one finds that the time required for the black hole mass to grow by a factor $X$ (and the black hole spin parameter to decrease from $\alpha_{0}$ to $\alpha=\alpha_{0}/X^{2+2\delta}$) is $t_{X}=(1-1/X)c^{5}/(G^{2}M_{0}\mathcal{A}\varepsilon_{\infty})$, where for simplicity we have set $t_{0}=0$. Therefore, the characteristic time scale $t_{\infty}$ for having a significant mass growth is related to the parameter $z$ according to
@@
VI acknowledgments
We thank Emilio Tejeda for fruitful discussions. P. M. acknowledges a support of the Polish National Science Centre Grant No. 2017/26/A/ST2/00530. M. M. was supported by CONAHCyT through Estancias Posdoctorales por México Convocatoria 2023(1) under the postdoctoral Grant No. 1242413. O. S. was partially supported by CIC Grant No. 18315 to Universidad Michoacana and by CONAHCyT Network Project No. 376127 “Sombras, lentes y ondas gravitatorias generadas por objetos compactos astrofísicos”. Finally, M. M. and O. S. acknowledge financial support from SECIHTI-SNII.

### v3

@@ Abstract
Introduction. The search for an exact solution representing a Bondi-type accretion in Kerr spacetime is a long-standing challenge. In 1988 Petrich, Shapiro, and Teukolsky obtained an analytic solution, assuming a zero-vorticity potential flow and the ultra-hard equation of state [1]. Their method was based on the observation that for the ultra-hard equation of state, the equations of motion governing zero-vorticity perfect fluids can be expressed in terms of a potential, and they are linear. More precisely, the continuity equation can be written as the massless Klein-Gordon equation, which is separable on the Kerr background spacetime. Unfortunately, no straightforward generalizations to perfect fluids with more realistic equations of state seem to be possible (see [2, 3, 4] for attempts to obtain approximate solutions).
@@
By a Bondi-type accretion we mean a nearly spherically symmetric, stationary flow of gas onto the black hole. The gas extends to infinity, where it is assumed to be homogeneous and at rest. Within the general-relativistic kinetic approximation, such models were investigated in [5, 6, 7, 8, 9, 10, 11], assuming spherical symmetry. Properties of other kinetic models in the Kerr spacetime were analyzed in [12, 13, 14, 15, 16]. Numerical simulations of accretion flows of the collisionless magnetized gas onto Kerr black holes were performed in [17, 18, 19, 20].
@@
Kinetic description of matter in the Kerr spacetime. Let $(\mathcal{M},g)$ and $T_{x}^{\ast}\mathcal{M}$ denote the spacetime manifold and the cotangent space at $x\in\mathcal{M}$, respectively. The gas is described in terms of the one-particle distribution function $f\colon\Gamma_{m}^{+}\to\mathbb{R}$ defined on the future mass shell $\Gamma_{m}^{+}:=\bigcup_{x\in\mathcal{M}}\{(x,p)\colon p\in P_{x}^{+}(m)\}$, where
@@
$p\text{ is future directed}\}.$
@@
Geodesic equations are separable [22, 23] in the coordinates $(t,r,\vartheta,\varphi)$ as follows: $p_{t}=-E$, $p_{\varphi}=L_{z}$, $p_{\vartheta}^{2}=\Theta(\vartheta)$, $\left(\Delta p_{r}-2MEr+aL_{z}\right)^{2}=R(r)$, where $\Delta:=r^{2}-2Mr+a^{2}$,
@@
Bondi-type solution. The components of the particle current density associated with absorbed and scattered orbits can be expressed as
@@
$\frac{\mathcal{\dot{J}}}{\varepsilon_{\infty}}$ $=$ $2\delta\mathcal{A}(1+\gamma\alpha^{2})M^{3}\alpha,$ (20b)
where
@@
which, using the relation $J=\alpha M^{2}$, implies a decrease of the spin parameter’s magnitude at the rate
$\frac{d\alpha}{dt}=-2\varepsilon_{\infty}\mathcal{A}(1+\delta)(1-\mathcal{C}\alpha^{2})M\alpha,\quad\mathcal{C}=\frac{\mathcal{B}-\delta\gamma}{1+\delta}.$ (21c)
Equations (21a) and (21c) govern the time evolution of the black hole mass and spin parameter, and they yield the following relation:
@@
The second scenario corresponds to a supermassive black hole which accretes dark matter from its environment. Assuming a reservoir of constant density and low temperature, neglecting contributions from baryonic matter which does not interact directly with the dark matter, and integrating Eq. (23), one finds that the time required for the black hole mass to grow by a factor $X$ (and the black hole spin parameter to decrease from $\alpha_{0}$ to $\alpha=\alpha_{0}/X^{2+2\delta}$) is $t_{X}=(1-1/X)c^{5}/(G^{2}M_{0}\mathcal{A}\varepsilon_{\infty})$, where for simplicity we have set $t_{0}=0$. Therefore, the characteristic time scale $t_{\infty}$ for having a significant mass growth is related to the parameter $z$ according to
@@
Acknowledgments. We thank Emilio Tejeda for fruitful discussions. P. M. acknowledges a support of the Polish National Science Centre Grant No. 2017/26/A/ST2/00530. M. M. was supported by SECIHTI through Estancias Posdoctorales por México Convocatoria 2023(1) under the postdoctoral Grant No. 1242413. O. S. was partially supported by CIC Grant No. 18315 to Universidad Michoacana and by CONAHCyT Network Project No. 376127 “Sombras, lentes y ondas gravitatorias generadas por objetos compactos astrofísicos”. Finally, M. M. and O. S. acknowledge financial support from SECIHTI-SNII.
Data availability. The data that support the findings of this article are openly available [34].